Context What follows is a data table showing the development of brilliance among a small class of PHE 450 students. NOTE: Student #8 came in as an existing case of brilliance and did not develop brilliance as a result of exposure to PHE 450. StudentWK 1 WK 2 WK 3 WK 4 WK 5 WK 6 WK 7 WK 8 WK 9 WK 10 1 2 3 4 CASE CASE CASE CASE DROP 7 8 CASE 9 10 CASE CASE DROP CASE 12 DROP1) what is the point of prevalence at brilliance at the end of week 1

2) what is the point of prevalence at brilliance at the end of week 2

3) what is the point of prevalence at brilliance at the end of week 3

4) Using person-weeks as your denominator, what is the incidence of brilliance over the course of the 10-week course? Provide your answer per 10 person-weeks.

5) Pretend the observation period of study was only the first three weeks. What is the cumulative incidence of brilliance at the end of three weeks of observation? Provide the cumulative incidence per 100 students.
Answer #1

Answer Prevalence =
no. of cases/population size
a. Prevalence can be measured in an closed cohort or in an open population.
b. Prevalence in cross-sectional.
c. “Old” cases and “new” cases are counted in the numerator.
d. Can be measured at a particular point (point prevalence) or over a period (period prevalence).

1.Point of prevalence at brilliance at the end of week 1= No.of cases/ population size.

no.of cases=

student 2 = 7 ( case of brilliance)

student3= 8 ( case of brilliance)

student 4= 2 (case of brilliance)

student 5=4 ( case of brilliance)

like wise total case of brilliance as per chart= 51 case of brilliance

population size=450

point of prevalence= 51/450= 11.33%

Situation2 :- Point of prevalence at end of week 2

point of prevalence= 51/450=11.33%

situation 3:- Point of prevalence at end of week 3

point of prevalence=49/450=10.88

no.of cases=49

population size 450.

situation 4.Incidence proportion = Cumulative Incidence = Risk =
no. of brilliance onsets/no. initially at risk
a. Incidence proportion can only be measured in a closed cohort.
b. Only new onsets are considered.
c. The time of follow-up must be specified
using person as denominater , incidence of brilliance=51/375=13.6